Situation: Joe, a DOH COVID-19 tracker, found out that the number of infected persons of COVID-19 in seven low-risk provinces of the country are 1, 9, 2, 12, 8, 7, and 10. Suppose that 5 provinces are drawn as sample.


1. What is the mean (u), variance (²) and standard deviation of the population (o)? 2. How many different samples of size n-5 can be drawn from the population?


List them with their corresponding mean. 3. Construct the sampling distribution of the sample mean.


4. What is the mean () of the sampling distribution of the sample mean? Compare this to the mean of the population.


5. What is the variance of the (o,) of the sampling distribution of the sample mean?


1
Expert's answer
2021-06-22T14:21:08-0400

1. We have population values "1,2,7,8,9,10,12" population size "N=7" and sample size "n=5."

Mean


"\mu=\dfrac{1+2+7+8+9+10+12}{7}=7"


Variance


"\sigma^2=\dfrac{1}{7}\big((1-7)^2+(2-7)^2+(7-7)^2"




"(8-7)^2+(9-7)^2+(10-7)^2+(12-7)^2\big)"




"=\dfrac{100}{7}"

Standard deviation


"\sigma=\sqrt{\sigma^2}=\sqrt{\dfrac{100}{7}}\approx3.779645"



2. The number of possible samples which can be drawn without replacement is


"\dbinom{N}{n}=\dbinom{7}{5}=21""\def\arraystretch{1.5}
 \begin{array}{c:c:c}
 Sample & Sample & Sample \ mean \\
 No. & values & (\bar{X}) \\ \hline
 1 & 1,2, 7, 8, 9 & 5.4 \\
 \hdashline
 2 & 1,2, 7, 8, 10 & 5.6 \\
 \hdashline
 3 & 1,2, 7,8,12 & 6.0 \\
 \hdashline
 4 & 1,2, 7, 9, 10 & 5.8\\
 \hdashline
 5 & 1,2,7,9,12 & 6.2 \\
\hdashline
 6 & 1,2, 7, 10,12 & 6.4 \\
 \hline
 7 & 1,2,8,9,10 & 6.0 \\
 \hline
 8 & 1,2,8,9,12 & 6.4\\
 \hline
 9 & 1,2,8,10,12 & 6.6 \\
 \hline
 10 & 1,2,9,10,12 & 6.8 \\
 \hline
 11 & 1,7,8,9,10 & 7.0 \\
 \hline
12 & 1,7,8,9,12 & 7.4 \\
 \hline
13 & 1,7, 8,10,12 & 7.6 \\
 \hline
14 & 1,7,9,10,12 & 7.8 \\
 \hline
15 & 1,8,9,10,12& 8.0 \\
 \hline
16 & 2,7,8,9,10& 7.2 \\
 \hline
17 & 2,7,8,9,12 & 7.6 \\
 \hline
18 & 2,7,8,10,12 & 7.8\\
 \hline
19 & 2,7,9,10,12 & 8.0 \\
 \hline
20 & 2,8,9,10,12 & 8.2 \\
 \hline
21 & 7,8, 9, 10, 12 & 9.2 \\
 \hline
 
\end{array}"



3.


"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f & f(\bar{X}) & \bar{X}f(\bar{X})& \bar{X}^2f(\bar{X}) \\ \hline
5.4 & 1 & 1/21 & 5.4/21 & 29.16/21 \\
 \hdashline
 5.6 & 1 & 1/21 & 5.6/21 & 31.36/21 \\
 \hdashline
 5.8 & 1 & 1/21& 5.8/21 & 33.64/21\\
 \hdashline
 6.0 & 2 & 2/21 & 12.0/21 & 72/21 \\
 \hdashline
 6.2 & 1 & 1/21 &6.2/21 & 38.44/21\\
 \hdashline
 6.4 & 2 & 2/21 & 12.8/21 & 81.92/21\\
 \hdashline
 6.6 & 1 & 1/21 & 6.6/21 & 43.56/21\\
 \hdashline
 6.8 & 1 & 1/21 & 6.8/21 & 46.24/21 \\
 \hdashline
 7.0 & 1 & 1/21 & 7.0/21 & 49/21\\
 \hdashline
 7.2 & 1 & 1/21 & 7.2/21 & 51.84/21 \\
 \hdashline
 7.4 & 1& 1/21 & 7.4/21 & 54.76/21\\
 \hdashline
 7.6 & 2 & 2/21 & 15.2/21 & 115.52/21\\
 \hdashline
 7.8 & 2 & 2/21 & 15.6/21 & 121.68/21\\
 \hdashline
 8.0 & 2 & 2/21 & 16.0/21 & 128/21\\
 \hdashline
 8.2 & 1& 1/21 & 8.2/21 & 67.24/21\\
 \hdashline
 9.2 & 1 & 1/21 & 9.2/21 & 84.64/21\\
 \hdashline
 Total & 21 & 1 & 7 & 1049/21 \\ \hline
\end{array}"




4.


"E(\bar{X})=\sum\bar{X}f(\bar{X})=7"

The mean of the sampling distribution of the sample means is equal to the

the mean of the population.



"E(\bar{X})=7=\mu"



5.


"Var(\bar{X})=\sum\bar{X}^2f(\bar{X})-(\sum\bar{X}f(\bar{X}))^2"




"=\dfrac{1049}{21}-(7)^2=\dfrac{20}{21}"




"\sqrt{Var(\bar{X})}=\sqrt{\dfrac{20}{21}}\approx0.975900"


Verification:


"Var(\bar{X})=\dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})=\dfrac{\dfrac{100}{7}}{5}(\dfrac{7-5}{7-1})"




"=\dfrac{20}{21}, True"